OpenStudy (anonymous):

What additional information would you need to prove the triangles congruent by ASA?

OpenStudy (anonymous):

What additional information would you need to prove the triangles congruent by ASA?

OpenStudy (anonymous):

A. SAS B. ASA C. AAS D. SSA

OpenStudy (anonymous):

omg please god bless all of you just help me with 10 questions

OpenStudy (nikato):

which one are you asking?

OpenStudy (anonymous):

for both i have 10 questions in all

OpenStudy (nikato):

1. what side do they share in common?

OpenStudy (anonymous):

fg and ih

OpenStudy (anonymous):

??

OpenStudy (paxpolaris):

right, FG = HI and angle F = angle H ... so to get angle-side-angle (ASA) .. you need the angle from the other end of the side to be congruent as well

OpenStudy (anonymous):

okay...and so now what

OpenStudy (anonymous):

so the answer is fi=gh

OpenStudy (anonymous):

that is one of the answers

OpenStudy (paxpolaris):

no that is already given in the graph. for ASA you 1 pair sides congruent , and the angles on either end of the sides congruent.

OpenStudy (anonymous):

so it would be g=i

OpenStudy (paxpolaris):

thats it :)

OpenStudy (anonymous):

yay okay so i have quesition 2 up there

OpenStudy (paxpolaris):

what are you thinking the answer to that is?

OpenStudy (anonymous):

sas

OpenStudy (paxpolaris):

for SAS, the congruent angle marked would have to be between the two pairs of congruent sides.

OpenStudy (anonymous):

oh so now what

OpenStudy (anonymous):

c?

OpenStudy (paxpolaris):

no 2 side have been marked and only one angle ... so 2 S and 1 A... but the angle is not in the middle of the sides so SSA

OpenStudy (anonymous):

oh okay let me get the next question sir

OpenStudy (paxpolaris):

note: SSA is a special case which only works when the congruent angles are right angles

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

how do i screenshot my screen?

OpenStudy (anonymous):

i want to show you because there all graphs

OpenStudy (paxpolaris):

press "Prt Scr" button on your keyboard ... then paste it in Paint

OpenStudy (anonymous):

find the length of ut and messure r

OpenStudy (anonymous):

okay ill try it

OpenStudy (paxpolaris):

wait ... for the other one you were write about SAS sorry

OpenStudy (anonymous):

awesome okay

OpenStudy (anonymous):

can you please just give me the answers god bless you cuz i have only 10 mins left its timed

OpenStudy (anonymous):

and i have like 7 more

OpenStudy (paxpolaris):

because SSA only applies to right angles it is not called that... it is called HS (hypotenuse-side)

OpenStudy (anonymous):

there is a question up there

OpenStudy (paxpolaris):

14x-36=2x+84 solve for x then plug it back in to get the length of the side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

10

OpenStudy (paxpolaris):

right plug it in 2x+84 to get UT

OpenStudy (anonymous):

b

OpenStudy (paxpolaris):

idk what b is. if that's the only one with UT =104 sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about r?

OpenStudy (paxpolaris):

R= U 10y-14=5y+56

OpenStudy (anonymous):

i got 126 for r

OpenStudy (paxpolaris):

5y=56+14 5y=70 y=14 yes 126

OpenStudy (anonymous):

okay i have more idk how to show u

OpenStudy (paxpolaris):

what computer are you using? windows7 / mac?

OpenStudy (anonymous):

window

OpenStudy (anonymous):

thanks im done i got a 60 god bless you and have a wonderful night

OpenStudy (paxpolaris):

good night